OpenStudy (toxicsugar22):

solve for r: A=P(1+r)^2

OpenStudy (anonymous):

A/P=(1+r)^2 +/- sqrt(A/P)=1+r r=sqrt(A/P)-1 or r=-sqrt(A/P)-1

OpenStudy (toxicsugar22):

thanks can u help me with one more

OpenStudy (toxicsugar22):

rationilixe the denominator 7a/sqrt6

OpenStudy (anonymous):

multiply top and bottom by sqrt6 7a*sqrt(6)/6 now the denominator is rational.

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

last one im sorry

OpenStudy (toxicsugar22):

nm i got that also

OpenStudy (toxicsugar22):

that is it thanks so much

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

looks like someone is doing school wrk ova here..lol